Entries for Investment in Bonds, Interest, and Sale of Bonds

The following bond investment transactions were completed during a recent year by Starks Company:

Year 1
Jan. 31 Purchased 45, $1,000 government bonds at 100 plus accrued interest of $225 (one month). The bonds pay 6% annual interest on July 1 and January 1.
July 1 Received semiannual interest on bond investment.
Aug. 30 Sold 18, $1,000 bonds at 97 plus $180 accrued interest (two months).

a. Journalize the entries for these transactions.

Assume a 360-day year. Do not round interim calculations. Round final answers to nearest dollar. For a compound transaction, if an amount box does not require an entry, leave it blank.

Year 1, Jan. 31
Year 1, July 1
Year 1, Aug. 30

b. Provide the December 31, Year 1, adjusting journal entry for semiannual interest earned on the bonds.

Assume a 360-day year. Do not round interim calculations. Round final answers to nearest dollar.

Year 1, Dec. 31
Answer #1

Answer:

Bonds purchased at 100 means they are purchased at par, below are the journal entries along with workings:

Date

Particulars

Debit

Credit

31-Jan

Investment in Bonds

   45000

Interest income

225

Cash account

   45,225

(To record purchase of 75 $1,000 bonds at 100 plus 30 days accrued interest)

Bonds amount = 45*1000 = 45,000

Interest amount = (45,000*6%)/12 = 225

1-Jul

Cash account

     1350

Interest income

     1350

(To record semi annual interest received)

Interest amount = (45,000*6%)/2 = 1350

29-Aug

Cash account

   17640

Loss on sale of bonds

         640

Interest income

         180

Investment in Bonds

   18000

(To record sale of 35 $1,000 bonds @98 plus interest earned $350)

Sale amount = 18*1000*97% = 17,460

Cash received = 17,460+180= 17,640

Bonds Original amount = 18*1000=18000

Loss on sale of bonds = 18000-17640= 640

31-Dec

Interest receivable account

     810

Interest income

     810

(To record semi annual interest earned on remaining 27 bonds)

Interest = (27*1000*6%)/2 = 810